OpenStudy (anonymous):

how do i turn the end points of a circle (1,1) and (-1,-1) in general form equation of a circle?

OpenStudy (amistre64):

end points of a circle? i assume you mean that is its diameter

myininaya (myininaya):

endpoints of a diameter of a circle?

OpenStudy (anonymous):

yes diameter

OpenStudy (amistre64):

find the midpoint; in this case it (0,0) which makes life easier

OpenStudy (anonymous):

yes

myininaya (myininaya):

midpoint of a diameter of a circle is the center of the circle

OpenStudy (amistre64):

and recall that the radius from (0,0) to (1,1) = sqrt(2)

myininaya (myininaya):

using distance formula to find the radius :)

OpenStudy (amistre64):

then fill in the circle equation: x^2 + y^2 = r^2

myininaya (myininaya):

if r=sqrt{2}, then r^2=?

OpenStudy (amistre64):

i think general form tho is what: x^2 + y^2 - r^2?

OpenStudy (anonymous):

\[(x-x _{1})(x-x _{2})+(y-y _{1})(y-y _{1})=0\]

OpenStudy (anonymous):

y-y2 in the last part

myininaya (myininaya):

x^2+y^2=r^2 is cool for centers not (0,0) use (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Substitute the values and you will eventually get the equation

myininaya (myininaya):

where (h,k) is center

OpenStudy (amistre64):

for centers not zero i throw my hands in the air and get some hagen daas

OpenStudy (anonymous):

how can u substitue 0s and get general form :/?

OpenStudy (amistre64):

(x-0)^2 = x^2

OpenStudy (amistre64):

same concept for the y part

myininaya (myininaya):

(h,k)=(0,0) then you have (x-0)^2+(y-0)^2=r^2 same as x^2+y^2=r^2

myininaya (myininaya):

ok later guys i have something to take care of

OpenStudy (amistre64):

to find the midpoint; i find it easier to add my points and divide it by 2 ( 1 , 1) (-1,-1) ------- ( 0 , 0 ) /2 = (0,0)

OpenStudy (anonymous):

same

OpenStudy (anonymous):

so then i also plug in 1,1 as x and y right?

OpenStudy (amistre64):

i call my center point: C(xy) to which (x-Cx)^2 + (y-Cy)^2 - r^2 = 0 would be the general form

OpenStudy (amistre64):

no; x and y are umm.. dummy variables that stand in for whatever it is we need them for

OpenStudy (amistre64):

Cx and Cy are what you should fill in

OpenStudy (anonymous):

oh so X^2+Y^2=R^2

OpenStudy (anonymous):

this is the gen form ?

OpenStudy (amistre64):

well, i answered prematurely; that is the standard form i believe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well its simplified so isnt it gen ?

OpenStudy (amistre64):

terminology gets me half the time; but standard from resembles the pythag thrm; general form is long and drawn out

OpenStudy (amistre64):

in this case its rather short and sweet; x^2 + y^2 -2 = 0

OpenStudy (anonymous):

wow had got this answer but it looked short reason i got confused

OpenStudy (amistre64):

thats becasue its centered at 0 and the extra stuff disappears

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)